OpenStudy (anonymous):

Solve (x+8)(2x-1)>_ 0 show all working

OpenStudy (anonymous):

the zeros are fairly clear, \[x+8=0\iff x=-8\] and \[2x-1=0\iff x=\frac{1}{2}\]

OpenStudy (anonymous):

since \[y=(x+8)(2x-1)\] is a parabola that opens up, it is negative between the zeros and positive outside them 

OpenStudy (anonymous):

you want to know where it is greater than or equal to zero that will be two intervals, \(x\leq -8\) or \(x\geq \frac{1}{2}\) which you can write that way, or in interval notation as \[(-\infty, -8]\cup [\frac{1}{2},\infty)\]